# Electronic structure of Ruddlesden-Popper nickelates: strain to mimic the effects pressure

### v1

@@ III.2 Trilayer RP nickelate
Moving on to the effects of strain in the electronic structure of La4Ni3O10, Fig. 4 shows the band structure plots (resolved for each inequivalent Ni atom in the trilayer structure: inner and outer) and the corresponding Fermi surfaces under a 3% compressive and tensile strain, in comparison to those obtained at 30 GPa. As shown in previous work [40, 47], the electronic structure of the trilayer RP at 30 GPa near the Fermi level is once again characterized by Ni-$e_{g}$ states hybridized with O-$p$ states. This is expected from formal valence considerations that in La4Ni3O10 give rise to an average $d^{7.33}$ filling (corresponding to a Ni2.67+ ion). As such, within each trilayer complex, four electrons need to be accommodated in the Ni-eg orbitals. As we have shown before [47], one should note that the $d_{z^{2}}$ orbitals are filled quite differently when comparing the inner-outer Ni atoms due to the formation of a bonding-nonbonding-antibonding molecular orbital complex with band splittings that are similar to those of La3Ni2O7 (the odd symmetry nonbonding state does not involve the inner Ni site as can be clearly observed in the fat bands). In contrast, no relevant distinction is obtained in the filling of the $d_{x^{2}-y^{2}}$ orbitals for the two inequivalent Ni atoms whose bandwidth is similar to that of the bilayer material at 30 GPa. Overall, five bands cross the Fermi level in La4Ni3O10 under pressure, consequently contributing to the Fermi surface. Following an equivalent notation to that used for the bilayer nickelate, we label these bands as $\delta$, $\alpha$, $\beta_{1}$, $\beta_{2}$, and $\gamma$. Similarly, the $\gamma$ pocket is purely $d_{z^{2}}$ in character (the small $\delta$ pocket is also $d_{z^{2}}$ dominated) while the $\alpha$, $\beta_{1}$, and $\beta_{2}$ sheets have a mixed $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbital character. The emergence of the $\gamma$ pocket is once again the main distinctive feature in the Fermi surface of La4Ni3O10 under high pressure when compared to the ambient pressure results and its appearance has been linked to the presence of an $s^{\pm}$ leading pairing symmetry - the calculated pairing strength of the $s^{\pm}$ gap structure being smaller than that obtained for the bilayer RP [40].
@@ III.3 Higher-order RPs
After having analyzed the effects of strain in the two nickelate RPs wherein superconducting signatures under pressure have been reported in the bulk, we now proceed to analyze the effects of strain in higher-order ($n>3$) RPs. As mentioned above, any RP nickelate with $n>3$ is not thermodynamically stable in bulk form but they can be synthesized in thin film form using molecular beam epitaxy [43, 44, 45]. These higher-order RPs could in principle enable to expand the family of RP nickelate superconductors, particularly if (epitaxial) strain could be used to mimic the effects of pressure.
@@
Given that the structures of the higher-order RP nickelates have not been experimentally resolved, we analyze the trends in the electronic structure under strain for the $n=4$ and 5 RPs (La5Ni4O13 and La6Ni5O16) using the corresponding tetragonal phases as a starting point at ambient pressure. We note that for both La3Ni2O7 and La4Ni3O10 the electronic structure in the tetragonalized phases at ambient pressure is very close to that of the orthorhombic or monoclinic cells [29, 37]. As applying large pressures to thin films will likely not be feasible, we simply compare the ambient pressure electronic structure of the $n=4$ and 5 RPs with that obtained under a 3% compressive and tensile strain to extract some general trends. As expected from their average electron fillings ($d^{7.25}$ for La5Ni4O13 and $d^{7.2}$ for La6Ni5O16), only bands of $e_{g}$ character cross the Fermi level, like in the bilayer and trilayer counterparts. As in the $n=2$ and 3 compounds, the filling of the $d_{x^{2}-y^{2}}$ bands does not significantly differ among the inequivalent Ni atoms in the structure while the filling of the $d_{z^{2}}$ orbitals is markedly different due to the formation of molecular orbitals (whose eigenvalues and eigenvectors are reported in previous work [29]). In the 4-layer RP La5Ni4O13, there are 5 bands crossing the Fermi level at ambient pressure: two smaller electron-pockets centered at $\Gamma$ ($\delta$ of pure $d_{z^{2}}$ character and $\alpha$ of mixed $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ character) and three larger sheets also of mixed orbital character: $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$. The electronic structure of the 5-layer nickelate is very similar, with the addition of one extra $\beta$-like sheet as a consequence of the extra Ni. For none of the higher-order RPs there are $\gamma$-like pockets at ambient pressure. Under compressive strain, these pockets are also absent and the fermiology looks similar to that at ambient pressure, other than for the splittings between the $\beta$-like pockets being largely increased at X. In contrast, under tensile strain the $\gamma$ pocket of pure $d_{z^{2}}$ character (associated with the emergence of $s^{\pm}$-wave pairing in the bilayer and trilayer compounds) is recovered, further confirming the trends we reported in the previous subsections. The same trends in the dominant hoppings described above for the bilayer and trilayer compounds are also obtained for the higher order RPs as shown in Fig. 9 of Appendix D.
@@
We have shown that strain is a viable way to decouple the structure and electronic-structure effects arising in Lan+1NinO3n+1 ($n=2-5$) RP nickelates under pressure. While tensile strain shortens the out-of-plane Ni-Ni distance, the Ni-O-Ni bond angle across the apical oxygen further deviates from 180 with respect to the ambient pressure cases. In contrast, compressive strain gives rise to larger out-of-plane Ni-Ni bond lengths and a Ni-O-Ni bond angle that approaches 180. In spite of the persistence of octahedral tilts, we have shown that tensile strain can mimic the effects of pressure in the electronic structure of the bilayer (La3Ni2O7) and trilayer (La4Ni3O10) RP nickelates, where superconducting signatures have been reported in bulk pressurized samples. In particular, tensile strain gives rise to $d_{z^{2}}$ pockets at the zone corners in the Fermi surface, whose presence has been associated with the emergence of s±-wave pairing. The appearance of these pockets under tensile strain can be extended to the higher-order RP nickelates $n=4-5$ (La5Ni4O13 and La6Ni5O16), making them interesting material platforms to study in order to expand superconductivity in this family beyond the trilayer compound. The decoupling that strain seems to provide between the structural and electronic structure characteristics arising under pressure makes it an interesting playground to investigate the competition between different pairing symmetries if superconductivity can be realized.
Note added. During the completion of this work, a preprint [60] appeared reporting the structure and electronic structure of strained La3Ni2O7 on LaAlO3 and SrTiO3. On SrTiO3 (that provides $\sim$ 1% tensile strain) the octahedral tilts are retained and the electronic structure also resembles that obtained under pressure.
@@ Appendix A MLWFs for RP nickelates
Appendix B Densities of states for RP nickelates
@@
Appendix C Electronic structure of 1% strained bilayer nickelate
@@
Appendix D Hopping parameters for $n=4-5$ RPs under strain

### v4

@@ III.2 Trilayer RP nickelate
Moving on to the effects of strain in the electronic structure of La4Ni3O10, Fig. 4 shows the band structure plots (resolved for each inequivalent Ni atom in the trilayer structure: inner and outer) and the corresponding Fermi surfaces under a 3% compressive and tensile strain, in comparison to those obtained at 30 GPa. As shown in previous work [40, 48], the electronic structure of the trilayer RP at 30 GPa near the Fermi level is once again characterized by Ni-$e_{g}$ states hybridized with O-$p$ states. This is expected from formal valence considerations that in La4Ni3O10 give rise to an average $d^{7.33}$ filling (corresponding to a Ni2.67+ ion). As such, within each trilayer complex, four electrons need to be accommodated in the Ni-eg orbitals. As we have shown before [48], one should note that the $d_{z^{2}}$ orbitals are filled quite differently when comparing the inner-outer Ni atoms due to the formation of a bonding-nonbonding-antibonding molecular orbital complex with band splittings that are similar to those of La3Ni2O7 (the odd symmetry nonbonding state does not involve the inner Ni site as can be clearly observed in the fat bands). In contrast, no relevant distinction is obtained in the filling of the $d_{x^{2}-y^{2}}$ orbitals for the two inequivalent Ni atoms whose bandwidth is similar to that of the bilayer material at 30 GPa. Overall, five bands cross the Fermi level in La4Ni3O10 under pressure, consequently contributing to the Fermi surface. Following an equivalent notation to that used for the bilayer nickelate, we label these bands as $\delta$, $\alpha$, $\beta_{1}$, $\beta_{2}$, and $\gamma$. Similarly, the $\gamma$ pocket is purely $d_{z^{2}}$ in character (the small $\delta$ pocket is also $d_{z^{2}}$ dominated) while the $\alpha$, $\beta_{1}$, and $\beta_{2}$ sheets have a mixed $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbital character. The emergence of the $\gamma$ pocket is once again the main distinctive feature in the Fermi surface of La4Ni3O10 under high pressure when compared to the ambient pressure results and its appearance has been linked to the presence of an $s^{\pm}$ leading pairing instability - the calculated pairing strength of the $s^{\pm}$ gap structure being smaller than that obtained for the bilayer RP [40].
@@ III.3 Higher-order RPs
After having analyzed the effects of strain in the two nickelate RPs wherein superconducting signatures under pressure have been reported in the bulk, we now proceed to analyze the effects of strain in higher-order ($n>3$) RPs. As mentioned above, any RP nickelate with $n>3$ is not thermodynamically stable in bulk form but they can be synthesized in thin film form using molecular beam epitaxy [44, 45, 46]. These higher-order RPs could, in principle, enable expanding the family of RP nickelate superconductors.
@@
Given that the structures of the higher-order RP nickelates have not been experimentally resolved, we analyze the trends in the electronic structure under strain for the $n=4$ and 5 RPs (La5Ni4O13 and La6Ni5O16) using the corresponding tetragonal phases as a starting point at ambient pressure. We note that for both La3Ni2O7 and La4Ni3O10 the electronic structure in the tetragonalized phases at ambient pressure is very close to that of the orthorhombic or monoclinic cells [29, 37]. As applying large pressures to thin films will likely not be feasible, we simply compare the ambient pressure electronic structure of the $n=4$ and 5 RPs with that obtained under a 3% compressive and tensile strain to extract some general trends. As expected from their average electron fillings ($d^{7.25}$ for La5Ni4O13 and $d^{7.2}$ for La6Ni5O16), only bands of $e_{g}$ character cross the Fermi level, like in the bilayer and trilayer counterparts. As in the $n=2$ and 3 compounds, the filling of the $d_{x^{2}-y^{2}}$ bands does not significantly differ among the inequivalent Ni atoms in the structure while the filling of the $d_{z^{2}}$ orbitals is markedly different due to the formation of molecular orbitals (whose eigenvalues and eigenvectors are reported in previous work [29]). In the 4-layer RP La5Ni4O13, there are 5 bands crossing the Fermi level at ambient pressure: two smaller electron-pockets centered at $\Gamma$ ($\delta$ of pure $d_{z^{2}}$ character and $\alpha$ of mixed $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ character) and three larger sheets also of mixed orbital character: $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$. The electronic structure of the 5-layer nickelate is very similar, with the addition of one extra $\beta$-like sheet as a consequence of the extra Ni. For none of the higher-order RPs there are $\gamma$-like pockets at ambient pressure. Under compressive strain, these pockets are also absent and the fermiology looks similar to that at ambient pressure, other than for the splittings between the $\beta$-like pockets being largely increased at X. In contrast, under tensile strain the $\gamma$ pocket of pure $d_{z^{2}}$ character (associated with the emergence of $s^{\pm}$-wave pairing in the bilayer and trilayer compounds) is recovered, further confirming the trends we reported in the previous subsections. The same trends in the dominant hoppings described above for the bilayer and trilayer compounds are also obtained for the higher order RPs as shown in Fig. 10 of Appendix E.
@@
By means of first-principles calculations, we have shown that strain is a viable way to decouple the structure and electronic-structure effects arising in Lan+1NinO3n+1 ($n=2-5$) RP nickelates under pressure. Compressive strain increases the out-of-plane Ni-Ni distance as well as the apical Ni-O-Ni angle bringing it closer to 180. In terms of the electronic structure, it shifts the $d_{z^{2}}$ flat bands away from the Fermi level giving rise to a more cuprate-like electronic structure. Such a shift of the $d_{z^{2}}$ flat bands under compressive strain may favor $d$-wave superconductivity. In contrast, tensile strain shortens the out-of-plane Ni-Ni distance and it decreases the apical Ni-O-Ni bond that further deviates from 180 with respect to the ambient pressure cases. In spite of the persistence of octahedral tilts, tensile strain recovers the flat $d_{z^{2}}$ bands crossing the Fermi level that have been associated with the emergence of $s^{\pm}$-wave pairing in the bilayer (La3Ni2O7) and trilayer (La4Ni3O10) RP nickelates. In this manner (biaxial) strain should be an interesting playground to pursue superconductivity in thin films of RP nickelates at ambient pressure as well as to investigate the competition between different pairing symmetries. Ultimately, it should also enable to pinpoint the structural and electronic structure characteristics necessary for superconductivity to emerge in this family of materials.
Note added. During the completion of this work, a preprint [61] appeared reporting the structure and electronic structure of strained La3Ni2O7 on LaAlO3 and SrTiO3. On SrTiO3 (that provides $\sim$ 1% tensile strain) the octahedral tilts are retained and the electronic structure resembles that obtained under pressure.
@@ Appendix A MLWFs for RP nickelates
Appendix B Electronic structure of strained bilayer nickelate La3Ni2O7 in $Amam$ symmetry
The band structures for La3Ni2O7 under a $\pm$3strain using an $Amam$ structure are shown in Fig. 7. The same trends in terms of band shifts as those observed in $Fmmm$ symmetry (described in the main text) are retained: note in particular the bonding $d_{z^{2}}$ band being shifted below the Fermi level (at M) under compressive strain while it clearly crosses it under tensile strain.
Figure 7: Band structure of La3Ni2O7 ($n=2$ RP) under different strains with orbital characters highlighted in $Amam$ symmetry: (a) for a $-3\%$ (compressive) strain, (b) for a $+3\%$ (tensile) strain.
Appendix C Densities of states for RP nickelates
@@
Appendix D Electronic structure of 1% strained bilayer nickelate
@@
Appendix E Hopping parameters for $n=4-5$ RPs under strain